# The NANOGrav 15 year Data Set: Removing pulsars one by one from the pulsar timing array

### v1

@@ Abstract
Evidence has emerged for a stochastic signal correlated among 67 pulsars within the 15-year pulsar-timing data set compiled by the NANOGrav collaboration. Similar signals have been found in data from the European, Indian, Parkes, and Chinese PTAs. This signal has been interpreted as indicative of the presence of a nanohertz stochastic gravitational wave background. To explore the internal consistency of this result we investigate how the recovered signal strength changes as we remove the pulsars one by one from the data set. We calculate the signal strength using the (noise-marginalized) optimal statistic, a frequentist metric designed to measure correlated excess power in the residuals of the arrival times of the radio pulses. We identify several features emerging from this analysis that were initially unexpected. The significance of these features, however, can only be assessed by comparing the real data to synthetic data sets. After conducting identical analyses on simulated data sets, we do not find anything inconsistent with the presence of a stochastic gravitational wave background in the NANOGrav 15-year data. The methodologies developed here can offer additional tools for application to future, more sensitive data sets. While this analysis provides an internal consistency check of the NANOGrav results, it does not eliminate the necessity for additional investigations that could identify potential systematics or uncover unmodeled physical phenomena in the data.
@@ 1 Introduction
[…]
Recently, several PTA collaborations have reported the results of the analyses of their most recent data sets showing the first emerging evidence for a spatially correlated process in the data. These results came from analysing the NANOGrav 15-year data set (NG15; Agazie et al., 2023a), the second data release from the European PTA (EPTA2; EPTA Collaboration et al., 2023) which includes the first data release from the Indian PTA (InPTA1; Joshi et al., 2018; EPTA Collaboration et al., 2023), the third data release from the Parkes PTA (PPTA3; Reardon et al., 2023) and the first data release from the Chinese PTA (CPTA1; Xu et al., 2023). In the near future, increased sensitivity is expected to be achieved using the forthcoming third data release from the International PTA (IPTA3) which will combine many of the aforementioned data sets, data from new observatories such as MeerKAT (which has already made its first data release, MPTA1; Miles et al., 2023) and eventually the Square Kilometre Array (SKA; Janssen et al., 2015). Despite the growing evidence for a stochastic gravitational-wave background (SGWB) there are good reasons to be cautious: tension with previous upper limits (Reardon et al., 2023); hints of unexpected angular correlation in some data sets (EPTA and NG15 monopole; EPTA Collaboration et al., 2023; Agazie et al., 2023a); differing signal characteristics between the published data sets and even from different time slices within single data sets (EPTA Collaboration et al., 2023; Reardon et al., 2023). Given the importance of any detection claim, it is necessary to verify the internal consistency of any data modeling (e.g., Agazie et al., 2024), especially for data sets that we intend to combine with others in further analysis. The objective of this paper is to devise tools to carry out such internal consistency checks. This work does not, however, negate the need for additional studies aimed at detecting potential systematics and identifying any unmodeled physical phenomena in the data.
In this paper we investigate the NANOGrav 15-year data set (NG15 hereafter). This data set is the product of a 67-pulsar PTA (those with a timing baseline $\geq$ 3 years), the most pulsars in any published PTA search currently. In the future, we plan to apply the methods developed here to other data sets, including those under construction within the IPTA.
@@ 2 Executive Summary
The NANOGrav collaboration recently published evidence for a SGWB in the 15-year data set (NG15; Agazie et al., 2023a). To deepen our understanding of the findings, we have developed an analysis pipeline that calculates how the recovered signal strength changes as pulsars are cumulatively removed from the PTA. Because our analysis requires the signal strength to be calculated thousands of times, we employ the noise marginalized optimal statistic, a computationally inexpensive frequentist detection metric, in our assessment of any signal. As well as a signal with the HD spatial correlations expected of a SGWB, we also searched for signals with monopole and dipole spatial correlations. A priori we anticipated that the recovered signal strength would exhibit a somewhat gradual decline as pulsars were removed from the array. However, the actual results show abrupt drops and even increases that were not anticipated. In order to quantify how unusual these features were, we designed metrics to quantify them and then compared the data to 100 realizations of simulated PTAs with a SGWB injected and with noise properties based on NG15. We find that the results of our pulsar-removal analysis of NG15 are consistent with the simulated data sets and, therefore, with a SGWB signal. It is still crucial, however, to pursue further studies that could expose underlying systematics or unmodeled physical influences in the data.
@@ 3 Methods
When using PTAs to search for a SGWB, multiple timing noise sources vary over timescales similar to the GW signal (e.g., intrinsic pulsar noise, interstellar-medium–induced radio-frequency-dependent fluctuations, and timing-model errors). To disentangle signal from noise, a joint estimate of both can be calculated using a Bayesian framework. By performing a fully Bayesian GW search, the analysis considers the relationships between parameters, provides a more nuanced understanding of the underlying astrophysical processes and incorporates all sources of uncertainty. However, a fully Bayesian analysis that incorporates HD spatial correlations requires substantial computing resources. Therefore, a useful complement to the Bayesian framework is the computationally inexpensive frequentist detection metric known as the optimal statistic (OS; Anholm et al., 2009).
@@ 3.1 The Optimal Statistic
The vector of timing residuals in each pulsar is denoted $\delta\mathbf{t}_{a}$, where the index $a$ labels the different pulsars in the array. When looking for a stochastic signal process that is spatially-correlated between different pulsars, the cross-covariance between two pulsars ($a\neq b$) is used;
@@
where $\sigma_{ab}$ are the uncertainties in the cross-correlations and where $\Gamma_{ab}$ is the overlap reduction function (ORF; quantifies the degree of correlation between pulsars as a function of their positions on the sky.) Under the assumptions that the noise is Gaussian and is much greater than the signal it can be shown that this statistic is optimal in the sense that it maximises the signal-to-noise ratio (S/N). Note that the autocorrelation terms involving $\rho_{aa}$ are intentionally omitted from the summation in this definition.
The OS defined above assumes a known ORF for a specific spatial correlation. However, in addition to the HD spatial correlations induced by a SGWB, other sources of spatially correlated timing residuals are possible. For example, clock errors can lead to monopole spatial correlations (Hobbs et al., 2012) and uncertainty in the position of the solar system barycenter can produce dipole spatial correlations (Champion et al., 2010) between the pulsars. The OS can be generalized to allow for a mixture of different possible spatial correlations (Sardesai et al., 2023). These multiple-component optimal statistic (MCOS) amplitudes, $A^{\alpha}$, are now the values that minimize
@@
$\Gamma^{\mathrm{Monopole}}_{ab}=1,\$ (4)
$\Gamma^{\mathrm{Dipole}}_{ab}=\cos\xi_{ab},\quad\mathrm{and}\$ (5)
$\begin{split}\Gamma^{\mathrm{HD}}_{ab}=\frac{\delta_{ab}}{2}+\frac{1}{2}-\frac{(1-\cos\xi_{ab})}{4}\left(\frac{1}{2}\right.\\ \left.-3\log\left[\frac{1-\cos\xi_{ab}}{2}\right]\right),\end{split}$ (6)
@@
In this paper, both the single-component (denoted OS; see Eq. 2) and multiple-component (denoted MCOS; see Eq. 3) are used. Single-component OS analyses are performed one-at-a-time using each of the three ORFs in Eqs. 4 to 6. Whenever the MCOS is used, it includes all three ORFs simultaneously.
The OS and the MCOS are computationally cheap to evaluate (for a typical PTA data set they can be evaluated in $\approx 0.5\;\mathrm{seconds})$. This characteristic makes them ideal for iterative analyses and facilitates the comprehensive exploration of a range of scenarios, such as different combinations of pulsars and spatial correlations and also when marginalizing over pulsar noise parameters (see Section 3.2).
3.2 The Noise-Marginalized Optimal Statistic
In cases where pulsars exhibit significant red noise, the OS yields biased results due to the strong covariance between individual red noise parameters and the amplitude of the SGWB. This bias can be mitigated by marginalizing over the individual pulsars’ red noise parameters using the posterior probability distributions from a Bayesian analysis of all PTA pulsars. This Bayesian step includes a SGWB red noise model that has a spectrum common to all pulsars, but which is agnostic regarding any spatial correlations between pulsars. This step must not be confused with a Bayesian analysis complete with spatial correlations, which is a much more computationally expensive process (Vigeland et al., 2018). In practise, each time the OS is calculated, the pulsars’ individual red noise parameters are drawn from the posterior distributions of the spatially uncorrelated Bayesian analysis; we calculate the OS 10,000 times, resulting in a distribution from which an average OS can be calculated. This method more accurately determines the strength of a signal embedded in the pulsar timing residuals than a single OS calculation and is many orders of magnitude faster than performing a Bayesian analysis complete with spatial correlations. We, therefore, use the noise-marginalized OS (NMOS) throughout this work.
@@ 3.2.1 Spatially Uncorrelated Bayesian Analysis
As described in Section 3.2, to calculate the NMOS we require the posterior probability distributions that result from a spatially uncorrelated Bayesian analysis of the pulsar data, which ignores spatial correlations, and we briefly describe that analysis here. When looking for the effects of gravitational waves, the pulsar data of interest are the timing residuals, which represent the deviations between observed and predicted pulse arrival times. For each pulsar we use the NG15 pulse times of arrival (TOAs) and the NG15 noise models
to compute the timing residuals. We assume that the timing residuals of the pulsars are produced by various physical processes characterized by variable parameters. Bayes’ theorem is then employed to deduce the probability distribution of these parameters of interest. Under the assumption of accurate noise modeling for each pulsar, the physical processes assumed to give rise to systematic trends seen in a pulsar’s timing residuals are:
@@
    The SGWB. The SGWB will imprint a common spectrum on all pulsar timing residuals. We model this SGWB spectrum as a $-2/3$ characteristic-strain power law, translating to a $-13/3$ power law for timing residuals. This power law is expected if the background is due to the incoherent superposition of radiation from SMBHBs in circular orbit in-spiraling due to loss of energy and angular momentum produced by GW emission  (Phinney, 2001). The amplitude of the SGWB spectrum is left as a free parameter and we assumed a log-uniform prior probability distribution between $10^{-18}$ and $10^{-11}$.
@@
    Individual pulsar red noise (also known as timing noise). Suggested interpretations of pulsar red noise include intrinsic changes in the spindown rate of the pulsar (Kaspi et al., 1994), interstellar propagation effects (Armstrong, 1984; Rickett, 1990; Kaspi et al., 1994; Cognard et al., 1995), and the presence of a circumpulsar asteroid belt (Shannon et al., 2013). The red noise for each pulsar is also modeled as a power law. The amplitude and the slope of the spectrum are both free parameters. We assumed a log-uniform prior probability distribution between $10^{-20}$ and $10^{-11}$ for the amplitude of the red noise and a uniform prior probability distribution between 0 and 7 for its slope.
@@
Each time a pulsar is removed from the PTA (as described in the next section) this Bayesian analysis step is completed anew with fewer pulsars. The red noise parameters of each pulsar subarray are then drawn from the Bayesian posterior distributions as described in Section 3.2.
3.3 Pulsar-removal analysis
Our aim is to use the above methods to quickly search for a correlated signal in the timing residuals of a pulsar timing array. Any signal detected will vary as a function of the number of pulsars in the array. One by one we eliminate pulsars from the full array and observe how the recovered signal is affected. The order in which we removed pulsars is based on their noise levels, specifically the weighted root-mean-square of epoch-averaged post-fit timing residuals after whitening (i.e., any red-noise contribution is removed) as listed in Table 6 of Agazie et al. (2023b). [^1] For PSR B1937+21, the value listed in Table 6 of Agazie et al. (2023b) is incorrect. This was discovered in the latter stages of this analysis. As a result, PSR B1937+21 appears at a position in the pulsar-removal ordering that reflects the value quoted in the Agazie et al. (2023b) table. We implemented two distinct exclusion methodologies for the NG15 pulsars: one where the least-noisy pulsars were excluded first, and another where the most-noisy pulsars were excluded first. A priori we anticipate that the removal of the least-noisy pulsars will have the most pronounced impact on the signals recovered from the timing residuals. It is for this reason that we prioritize the scenario in which the least-noisy pulsars are removed first. The case where we remove the most-noisy pulsars first is discussed in the appendix of this paper.

### v2

# The NANOGrav 15 Yr Data Set: Removing Pulsars One by One from the Pulsar Timing Array

@@ Abstract
Evidence has emerged for a stochastic signal correlated among 67 pulsars within the 15 yr pulsar-timing data set compiled by the NANOGrav collaboration. Similar signals have been found in data from the European, Indian, Parkes, and Chinese pulsar timing arrays. This signal has been interpreted as indicative of the presence of a nanohertz stochastic gravitational-wave background (GWB). To explore the internal consistency of this result, we investigate how the recovered signal strength changes as we remove the pulsars one by one from the data set. We calculate the signal strength using the (noise-marginalized) optimal statistic, a frequentist metric designed to measure the correlated excess power in the residuals of the arrival times of the radio pulses. We identify several features emerging from this analysis that were initially unexpected. The significance of these features, however, can only be assessed by comparing the real data to synthetic data sets. After conducting identical analyses on simulated data sets, we do not find anything inconsistent with the presence of a stochastic GWB in the NANOGrav 15 yr data. The methodologies developed here can offer additional tools for application to future, more sensitive data sets. While this analysis provides an internal consistency check of the NANOGrav results, it does not eliminate the necessity for additional investigations that could identify potential systematics or uncover unmodeled physical phenomena in the data.
@@ 1 Introduction
Pulsar timing arrays (PTAs) can be used to detect gravitational waves (GWs) in the nanohertz frequency band by examining the timing residuals in a network of millisecond pulsars (MSPs) (Sazhin, 1978; Detweiler, 1979; Foster & Backer, 1990). The timing residuals induced by a Gaussian and isotropic stochastic background of GWs in different pulsars are correlated across the sky in a characteristic way (described by the Hellings-Downs (HD) curve; Hellings & Downs, 1983) which allows them to be distinguished from other mismodelled processes. It is expected that inspiralling supermassive black holes binaries (SMBHB), with masses in the range $\sim$$\textsterling 10^{8}\,\textrm{--}\,10^{10}\,M_{\odot}$ and redshifts $z\lesssim 2$, will produce a stochastic GW background (SGWB) in this frequency band (Jaffe & Backer, 2003; Sesana et al., 2004, 2008; McWilliams et al., 2014), although individual sources and cosmological backgrounds may also be present with comparable amplitudes.
Recently, several PTA collaborations have reported the results of the analyses of their most recent data sets showing the first emerging evidence for a spatially correlated process in the data. These results came from analysing the NANOGrav 15 yr data set (NG15; Agazie et al., 2023a), the second data release from the European PTA (EPTA2; EPTA Collaboration et al., 2023) which includes the first data release from the Indian PTA (Joshi et al., 2018; EPTA Collaboration et al., 2023), the third data release from the Parkes PTA (Reardon et al., 2023) and the first data release from the Chinese PTA Xu et al. (2023). In the near future, increased sensitivity is expected to be achieved using the forthcoming third data release from the International PTA which will combine many of the aforementioned data sets, data from new observatories such as MeerKAT (which has already made its first data release, MPTA1; Miles et al., 2023) and eventually the Square Kilometre Array (SKA; Janssen et al., 2015). Despite the growing evidence for a stochastic gravitational-wave background (SGWB) there are good reasons to be cautious: tension with previous upper limits (Reardon et al., 2023); hints of unexpected angular correlation in some data sets (EPTA and NG15 monopole; EPTA Collaboration et al., 2023; Agazie et al., 2023a); differing signal characteristics between the published data sets and even from different time slices within single data sets (EPTA Collaboration et al., 2023; Reardon et al., 2023). Given the importance of any detection claim, it is necessary to verify the internal consistency of any data modeling (e.g., Agazie et al., 2024), especially for data sets that we intend to combine with others in further analysis. The objective of this paper is to devise tools to carry out such internal consistency checks. This work does not, however, negate the need for additional studies aimed at detecting potential systematics and identifying any unmodeled physical phenomena in the data.
In this paper we investigate the NANOGrav 15 yr data set (NG15 hereafter). This data set is the product of a 67-pulsar PTA (those with a timing baseline $\geq$ 3 years), the most pulsars in any published PTA search currently. In the future, we plan to apply the methods developed here to other data sets, including those under construction within the IPTA.
@@ 2 Executive Summary
The NANOGrav collaboration recently published evidence for an SGWB in the 15 yr data set (NG15; Agazie et al., 2023a). To deepen our understanding of the findings, we have developed an analysis pipeline that calculates how the recovered signal strength changes as pulsars are cumulatively removed from the PTA. Because our analysis requires the signal strength to be calculated thousands of times, we employ the noise marginalized (NM) optimal statistic (OS), a computationally inexpensive frequentist detection metric, in our assessment of any signal. As well as a signal with the HD spatial correlations expected of an SGWB, we also searched for signals with monopole and dipole spatial correlations. A priori we anticipated that the recovered signal strength would exhibit a somewhat gradual decline as pulsars were removed from the array. However, the actual results show abrupt drops and even increases that were not anticipated. In order to quantify how unusual these features were, we designed metrics to quantify them and then compared the data to 100 realizations of simulated PTAs with an SGWB injected and with noise properties based on NG15. We find that the results of our pulsar removal analysis of NG15 are consistent with the simulated data sets and, therefore, with an SGWB signal. It is still crucial, however, to pursue further studies that could expose underlying systematics or unmodeled physical influences in the data.
@@ 3 Methods
When using PTAs to search for an SGWB, multiple timing noise sources vary over timescales similar to the GW signal (e.g., intrinsic pulsar noise, interstellar-medium–induced radio-frequency-dependent fluctuations, and timing-model errors). To disentangle signal from noise, a joint estimate of both can be calculated using a Bayesian framework. By performing a fully Bayesian GW search, the analysis considers the relationships between parameters, provides a more nuanced understanding of the underlying astrophysical processes and incorporates all sources of uncertainty. However, a fully Bayesian analysis that incorporates HD spatial correlations requires substantial computing resources. Therefore, a useful complement to the Bayesian framework is the computationally inexpensive frequentist detection metric known as the OS (Anholm et al., 2009).
@@ 3.1 The Optimal Statistic
The vector of timing residuals in each pulsar is denoted $\delta\mathbf{t}_{a}$, where the index $a$ labels the different pulsars in the array. When looking for a stochastic signal process that is spatially correlated between different pulsars, the cross-covariance between two pulsars ($a\neq b$) is used;
@@
where $\sigma_{ab}$ are the uncertainties in the cross-correlations and where $\Gamma_{ab}$ is the overlap reduction function (ORF; which quantifies the degree of correlation between pulsars as a function of their positions on the sky.) Under the assumptions that the noise is Gaussian and is much greater than the signal it can be shown that this statistic is optimal in the sense that it maximizes the signal-to-noise ratio (S/N). Note that the autocorrelation terms involving $\rho_{aa}$ are intentionally omitted from the summation in this definition.
The OS defined above assumes a known ORF for a specific spatial correlation. However, in addition to the HD spatial correlations induced by an SGWB, other sources of spatially correlated timing residuals are possible. For example, clock errors can lead to monopole spatial correlations (Hobbs et al., 2012) and uncertainty in the position of the solar system barycenter can produce dipole spatial correlations (Champion et al., 2010) between the pulsars. The OS can be generalized to allow for a mixture of different possible spatial correlations (Sardesai et al., 2023). These multiple-component optimal statistic (MCOS) amplitudes, $A^{\alpha}$, are now the values that minimize
@@
$\Gamma^{\rm Monopole}_{ab}=1,$ (4)
$\Gamma^{\rm Dipole}_{ab}=\cos\xi_{ab},\quad\mathrm{and}$ (5)
$\Gamma^{\rm HD}_{ab}=\frac{\delta_{ab}}{2}+\frac{1}{2}-\frac{(1-\cos\xi_{ab})}{4}\bigg{(}\frac{1}{2}$
$\hskip 28.45274pt-3\log\left[\frac{1-\cos\xi_{ab}}{2}\right]\bigg{)},$ (6)
@@
In this paper, both the single component (SC) OS (see Eq. 2) and MCOS (see Eq. 3) are used. SCOS analyses are performed one at a time using each of the three ORFs in Eqs. 4 to 6. Whenever the MCOS is used, it includes all three ORFs simultaneously.
The OS and the MCOS are computationally cheap to evaluate (for a typical PTA data set they can be evaluated in $\approx 0.5\;\mathrm{s})$. This characteristic makes them ideal for iterative analyses and facilitates the comprehensive exploration of a range of scenarios, such as different combinations of pulsars and spatial correlations and also when marginalizing over pulsar-noise parameters (see Section 3.2).
3.2 The NMOS
In cases where pulsars exhibit significant red noise, the OS yields biased results due to the strong covariance between individual red-noise parameters and the amplitude of the SGWB. This bias can be mitigated by marginalizing over the individual pulsars’ red-noise parameters using the posterior probability distributions from a Bayesian analysis of all PTA pulsars. This Bayesian step includes an SGWB red-noise model that has a spectrum common to all pulsars, but that is agnostic regarding any spatial correlations between pulsars. This step must not be confused with a Bayesian analysis complete with spatial correlations, which is a much more computationally expensive process (Vigeland et al., 2018). In practice, each time the OS is calculated, the pulsars’ individual red-noise parameters are drawn from the posterior distributions of the spatially uncorrelated Bayesian analysis; we calculate the OS 10,000 times, resulting in a distribution from which an average OS can be calculated. This method more accurately determines the strength of a signal embedded in the pulsar timing residuals than a single OS calculation and is many orders of magnitude faster than performing a Bayesian analysis complete with spatial correlations. We, therefore, use the NMOS throughout this work.
@@ 3.2.1 Spatially Uncorrelated Bayesian Analysis
As described in Section 3.2, to calculate the NMOS we require the posterior probability distributions that result from a spatially uncorrelated Bayesian analysis of the pulsar data, which ignores spatial correlations, and we briefly describe that analysis here. When looking for the effects of gravitational waves, the pulsar data of interest are the timing residuals, which represent the deviations between observed and predicted pulse arrival times. For each pulsar we use the NG15 pulse times of arrival (TOAs) and the NG15 noise models to compute the timing residuals. We assume that the timing residuals of the pulsars are produced by various physical processes characterized by variable parameters. Bayes’ theorem is then employed to deduce the probability distribution of these parameters of interest. Under the assumption of accurate noise modeling for each pulsar, the physical processes assumed to give rise to systematic trends seen in a pulsar’s timing residuals are:
@@
    The SGWB. The SGWB will imprint a common spectrum on all pulsar timing residuals. We model this SGWB spectrum as a $-2/3$ characteristic-strain power law, translating to a $-13/3$ power law for timing residuals. This power law is expected if the background is due to the incoherent superposition of radiation from SMBHBs in circular orbit in-spiraling due to loss of energy and angular momentum produced by GW emission  (Phinney, 2001). The amplitude of the SGWB spectrum is left as a free parameter and we assume a log-uniform prior probability distribution between $10^{-18}$ and $10^{-11}$.
@@
    Individual pulsar red noise (also known as timing noise). Suggested interpretations of pulsar red noise include intrinsic changes in the spindown rate of the pulsar (Kaspi et al., 1994), interstellar propagation effects (Armstrong, 1984; Rickett, 1990; Kaspi et al., 1994; Cognard et al., 1995), and the presence of a circumpulsar asteroid belt (Shannon et al., 2013). The red noise for each pulsar is also modeled as a power law. The amplitude and the slope of the spectrum are both free parameters. We assume a log-uniform prior probability distribution between $10^{-20}$ and $10^{-11}$ for the amplitude of the red noise and a uniform prior probability distribution between 0 and 7 for its slope.
@@
Each time a pulsar is removed from the PTA (as described in the next section) this Bayesian analysis step is completed anew with fewer pulsars. The red-noise parameters of each pulsar subarray are then drawn from the Bayesian posterior distributions as described in Section 3.2.
3.3 Pulsar removal analysis
Our aim is to use the above methods to quickly search for a correlated signal in the timing residuals of a PTA. Any signal detected will vary as a function of the number of pulsars in the array. One by one we eliminate pulsars from the full array and observe how the recovered signal is affected. The order in which we removed pulsars is based on their noise levels, specifically the weighted root-mean-square of epoch-averaged post-fit timing residuals after whitening (i.e., any red-noise contribution is removed) as listed in Table 6 of Agazie et al. (2023b) [^1] For PSR B1937+21, the value listed in Table 6 of Agazie et al. (2023b) is incorrect. This was discovered in the latter stages of this analysis. As a result, PSR B1937+21 appears at a position in the pulsar removal ordering that reflects the value quoted in the Agazie et al. (2023b) table. We implemented two distinct exclusion methodologies for the NG15 pulsars: one where the least noisy pulsars were excluded first, and another where the most noisy pulsars were excluded first. A priori we anticipate that the removal of the least noisy pulsars will have the most pronounced impact on the signals recovered from the timing residuals. It is for this reason that we prioritize the scenario in which the least noisy pulsars are removed first. The case where we remove the most noisy pulsars first is discussed in Appendix A of this paper.